OpenStudy (anonymous):

Use implicit differentiation to find the equation of the tangent line to the curve at x^2 + xy + y^2 = 3 at (1,1) How does this problem start? i know we need to find the slope and b, etc right? but not sure how to work the problem... please explain! Thanks!

OpenStudy (phi):

The first step is use implicit differentiation to find dy/dx can you do that ?

OpenStudy (anonymous):

umm, do I do this? dy --- = chain rule? dx

OpenStudy (phi):

\[ \frac{d}{dx}\left( x^2 + xy + y^2 = 3\right) \\ \frac{d}{dx}x^2 + \frac{d}{dx}(xy) + \frac{d}{dx}y^2 = \frac{d}{dx}3 \\ \]

OpenStudy (phi):

\[ \frac{d}{dx}x^2 = 2x \frac{dx}{dx} = 2x \] \[ \frac{d}{dx}y^2 = 2y \frac{dy}{dx} \]

OpenStudy (phi):

the derivative of xy uses the product rule \[ \frac{d}{dx}(xy) = x \frac{d}{dx}y + y \frac{d}{dx}x =x \frac{dy}{dx} +y\] and of course d/dy 3 is zero solve for dy/dx (it is easier to first sub in 1 for x and y, because you want dy/dx at the point (1,1) ) to find the slope of the line tangent to the curve at that point. finally, use y - 1 = m(x - 1) for the equation of the line. m is dy/dx evaluated at (1,1)

OpenStudy (anonymous):

ohh okay... so m=0 in this problem? so y-1=0(x-1) ?

OpenStudy (phi):

\[ \frac{d}{dx}\left( x^2 + xy + y^2 = 3\right) \\ \frac{d}{dx}x^2 + \frac{d}{dx}(xy) + \frac{d}{dx}y^2 = \frac{d}{dx}3 \\ 2x + x \frac{dy}{dx} + y + 2 y \frac{dy}{dx}=0 \] treat dy/dx as a variable, and solve for dy/dx if we first sub in 1 for x and y (because we are evaluating this at (1,1) ) \[ 2 + \frac{dy}{dx}+1 +2\frac{dy}{dx}=0 \] what do you get for dy/dx ?

OpenStudy (anonymous):

umm 5 dy/dx=0 ? dy/dx=0 ?

OpenStudy (phi):

Have you reached the limits of your talents? how about rename dy/dx as X , so the equation is 2+X + 1 + 2X =0 collect like terms.

OpenStudy (anonymous):

ohh okay.. so 3+3x=0 ?

OpenStudy (anonymous):

3x=3? x=1?

OpenStudy (phi):

mucho bettero

OpenStudy (phi):

except I get -1

OpenStudy (anonymous):

oh right haha oops :P yes x=-1 :) so now we solve for the equation? y-1=-1(x-1) =y-1= -x +1 y= -x + 2 ?

OpenStudy (phi):

yes, that looks good.

OpenStudy (anonymous):

ah yay!! thank you!! :)

OpenStudy (phi):

If you are losing concentration on this stuff, you might take a breather. It takes concentration to remember and apply your knowledge.

OpenStudy (anonymous):

okay :) will do!! thank you1!!